OpenStudy (anonymous):

Solve this system algebraically. 7x - 2y = 4 5y + 3x = 10 {(-41/40, -41/58)} {(41/40, 41/58)} {(-40/41, -58/41)} {(40/41, 58/41)}

OpenStudy (anonymous):

21x - 6y = 12 21x + 35y = 70

OpenStudy (anonymous):

41y = 58 y = 58/41

OpenStudy (anonymous):

so it's {(40/41, 58/41)}

OpenStudy (anonymous):

y = 2x + 3 2y = 4x + 6 The system of equations has _____ solution(s). no one infinite

OpenStudy (anonymous):

infinite

OpenStudy (anonymous):

What is the x-coordinate of the solution of the following system of equations? 3x + y = 6 x - y = 6 -3 0 3

OpenStudy (anonymous):

3

OpenStudy (anonymous):

5x + 2y = 7 y = x + 1 What is the solution set of the given system? {(5/7, 12/7)} {(12/7, 5/7)} {(9/7, 16/7)} {(16/7, 9/7)}